OpenStudy (anonymous):

A traffic light is red for 38 seconds, yellow for 10 seconds, green for one minute and twelve seconds in a constantly repeating cycle. What is the probability that at the moment a passing pedestrian glances at it, the light will be green? The answer is 3/5 or .6 but the question is how to do the solution to that?

OpenStudy (anonymous):

First, find how many seconds the light is green: 60+12=72. Find the number of seconds it takes to go through one cycle: 38+10+72=120. Divide the seconds it's green by total : 72/120=(3*24)/(5*24)=3/5.

OpenStudy (anonymous):

Thank you so much.